OpenStudy (anonymous):

A carton contains 12 light bulbs of which 3 are defective. if a random sample of 4 light bulbs is chosen without replacement, what is the probability distribution of defective light bulbs?

OpenStudy (kropot72):

\[P(0\ defective)=\frac{3C0\times9C4}{12C4}\] \[P(1\ defective)=\frac{3C1\times9C3}{12C4}\] \[P(2\ defective)=\frac{3C2\times9C2}{12C4}\] \[P(3\ defective)=\frac{3C3\times9C1}{12C4}\]

OpenStudy (anonymous):

Ahhhhh!

OpenStudy (anonymous):

Thank u <3

OpenStudy (kropot72):

You're welcome :)